# A Node on the Constellation: The Role of Feminist Makerspaces in Building and Sustaining Alternative Cultures of Technology Production

### v1

@@ 5.1. The Necessity of Long-Term Stewardship
Finally, creators of countertechnologies, those building viable alternatives to extractive platforms, face what Tandon et al. (Tandon et al., 2022) describe as “hostile ecologies”: restricted access to funding, inequitable regulatory frameworks, diminished design agency, and community partners with limited bandwidth or divergent priorities. These dynamics are familiar to those working on decentralized gig worker platforms, such as projects that Irani and Silberman (Irani and Silberman, 2013), and Calacci (Calacci, 2022) and others have explored. We observed similar forms of hostility in our study. As Gene’s account of displacement highlights, feminist makerspaces must be able to navigate hostile urban ecologies where rising rents and gentrification threaten their physical survival. Internal dynamics, too, can undermine collective care—as seen in the ousting of Dana, a Black trans woman whose grassroots outreach was dismissed by a founding member in favor of more “professional” approaches. Building and maintaining counterinstitutional systems requires not only technical expertise but persistent organizing, trust-building, and institutional repair, all of which are forms of stewardship. In this context, stewardship is not merely about keeping servers online or meeting grant deliverables. It is about cultivating the conditions under which feminist tech infrastructures can survive, through leadership succession planning, distributed knowledge sharing, collective ownership models, and alliances with other institutions of care. The future of feminist makerspaces depends not only on the moment of their creation but also on our collective willingness to sustain them.
@@ 5.2. Solidarity over Scale, in Order to Build Counterspaces
Throughout our interviews, founders and members of feminist makerspaces returned to a central question: How do we sustain our work in ways that center care, reciprocity, and collective survival? Rather than viewing sustainability as an internal organizational challenge or a question of longevity alone, participants reframed it through an interconnected lens of mutual aid and solidarity. Drawing on Dean Spade’s (Spade, 2020) work, we understand mutual aid as the radical act of meeting people’s immediate needs through collective care and bottom-up organizing, rather than relying on institutional or philanthropic gatekeeping. As seen when members rallied to sustain the first author’s space during the COVID-19 pandemic, or in efforts to keep tenant rents below market rate to prevent displacement, feminist makerspaces often rely on collective care and reciprocal support rather than external gatekeepers. Feminist makerspaces, then, are not only sites of technical production or education—they are critical infrastructures for redistributing resources and building interdependence across communities.
@@
By centering mutual aid and solidarity, feminist makerspaces have begun to articulate a vision of sustainability grounded not in scale, institutional prestige, or corporate partnerships, but in the everyday work of showing up for one another (of Color Against Violence, 2017). This approach demands that makerspaces remain responsive to current social conditions, and that they engage in collaborative struggle with other social justice organizations as part of a broader fabric of resistance and care.

### v2

@@ 5.1. The Necessity of Long-Term Stewardship
Finally, creators of countertechnologies, those building viable alternatives to extractive platforms, face what Tandon et al. (Tandon et al., 2022) describe as “hostile ecologies”: restricted access to funding, inequitable regulatory frameworks, diminished design agency, and community partners with limited bandwidth or divergent priorities. These dynamics are familiar to those working on decentralized gig worker platforms, such as projects that Irani and Silberman (Irani and Silberman, 2013), and Calacci (Calacci, 2022) and others have explored. We observed similar forms of hostility in our study. As Gene’s account of displacement highlights, feminist makerspaces must be able to navigate hostile urban ecologies where rising rents and gentrification threaten their physical survival. Internal dynamics, too, can undermine collective care—as seen in the conflict with Dana, a Black trans woman whose grassroots outreach was dismissed by a founding member in favor of more “professional” approaches. Building and maintaining counterinstitutional systems requires not only technical expertise but persistent organizing, trust-building, and institutional repair, all of which are forms of stewardship. In this context, stewardship is not merely about keeping servers online or meeting grant deliverables. It is about cultivating the conditions under which feminist tech infrastructures can survive, through leadership succession planning, distributed knowledge sharing, collective ownership models, and alliances with other institutions of care. The future of feminist makerspaces depends not only on the moment of their creation but also on our collective willingness to sustain them.
@@ 5.2. Solidarity over Scale, in Order to Build Counterspaces
Throughout our interviews, founders and members of feminist makerspaces returned to a central question: How do we sustain our work in ways that center care, reciprocity, and collective survival? Rather than viewing sustainability as an internal organizational challenge or a question of longevity alone, participants reframed it through an interconnected lens of mutual aid and solidarity. Drawing on Dean Spade’s (Spade, 2020) work, we understand mutual aid as the radical act of meeting people’s immediate needs through collective care and bottom-up organizing, rather than relying on institutional or philanthropic gatekeeping. As seen in efforts to keep tenant rents below market rate to prevent displacement, feminist makerspaces often rely on collective care and reciprocal support rather than external gatekeepers. Feminist makerspaces, then, are not only sites of technical production or education—they are critical infrastructures for redistributing resources and building interdependence across communities.
@@
By centering mutual aid and solidarity, feminist makerspaces have begun to articulate a vision of sustainability grounded not in scale, institutional prestige, or corporate partnerships, but in the everyday work of showing up for one another (of Color Against Violence, 2017). This approach demands that feminist makerspaces remain responsive to current social conditions, and that they engage in collaborative struggle with other social justice organizations as part of a broader fabric of resistance and care.